# Resona: Improving Context Copying in Linear Recurrence Models with Retrieval

### v1

@@ 3.2 Training and Inference
When training Resona, an important consideration is the auto-regressive nature of the model that therefore requires a causal mask. Specific to our implementation, to maintain this during the parallel nature of training, we ensure that for the query at index $i$, only chunks that solely contain information prior to this position in the input are considered within the Chunk-and-Search process. This ensures that the mask ${\bm{M}}$ allows no information from future tokens to a the representation from the past.
@@ 4.2 Main Results
Model Compression ICR Noisy ICR Fuzzy ICR SC Memorization Average
Mamba 38.3 76.7 74.9 9.3 33.2 88.5 53.5
   w/ Resona 38.2 ($\downarrow$ 0.1) 99.9 ($\uparrow$ 23.2) 100.0 ($\uparrow$ 25.1) 63.4 ($\uparrow$ 54.1) 42.7 ($\uparrow$ 9.5) 88.8 ($\uparrow$ 0.3) 72.1 ($\uparrow$ 18.6)
Mamba2 43.6 96.4 96.7 21.1 93.3 86.9 73.0
   w/ Resona 46.6 ($\uparrow$ 3.0) 100.0 ($\uparrow$ 3.6) 100.0 ($\uparrow$ 3.3) 62.9 ($\uparrow$ 41.8) 93.6 ($\uparrow$ 0.3) 88.1 ($\uparrow$ 1.2) 81.9 ($\uparrow$ 8.9)
RWKV5 36.8 96.4 96.6 12.1 52.7 55.0 58.3
   w/ Resona 40.4 ($\uparrow$ 3.6) 99.7 ($\uparrow$ 3.3) 99.8 ($\uparrow$ 3.2) 59.7 ($\uparrow$ 47.6) 58.0 ($\uparrow$ 5.3) 70.6 ($\uparrow$ 15.6) 71.5 ($\uparrow$ 13.2)
Hyena 42.2 79.3 77.4 9.96 72.8 88.9 61.7
   w/ Resona 42.6 ($\uparrow$ 0.4) 99.9 ($\uparrow$ 20.6) 99.9 ($\uparrow$ 22.5) 66.2 ($\uparrow$ 56.2) 74.3 ($\uparrow$ 1.5) 89.0 ($\uparrow$ 0.1) 78.7 ($\uparrow$ 17.0)
@@
DeltaNet (same param) 142M 12.946
DeltaNet+Resona 145M 12.541
@@
RetNet (same param) 146M 15.431
RetNet$+$Resona 142M 14.742
@@
Mamba (same param) 157M 16.173
Mamba$+$Resona 154M 15.943
@@ 4.2.2 Language Modeling
To assess language modeling, we compare a baseline models with one augmented with Resona layers. Here, we train on the WikiText-103 dataset (Merity et al., 2016). In order to train models augmented with Resona, we also augment the dataset. Specifically, we first consolidate all samples from the same Wikipedia entry into single sample, eliminating excessively short title lines or empty lines. We then use a LLaMA3.1-70B (Grattafiori et al., 2024) model to augment each sample such that we can make use of the Resona retrieval mechanism. We then conduct the Chunk-and-Search process offline to create masks, in order to save computation during training. To account for the additional parameters introduced by Resona, we present results in the baseline settings for two version, one where each LRM layer matches exactly that of the augmented model as well as a version where the hidden size of the layer has been increased to match the parameter count of the augmented counterpart. Evaluation results in Table 2 demonstrate the effectiveness of integrating Resona consistently achieve lower perplexity than both their baseline counterparts and parameter-aligned variants, highlighting the general applicability of Resona for language modeling. Notably, in the last row, modifying Hymba’s sliding window mask to our retrieve-lead mask significantly improves performance, further underscoring its advantage over fixed sliding window masking.
@@ 4.2.3 Direct Fine-Tuning
   w/ Resona 25.6 61.0 43.4 41.5 36.5 69.0 48.8 57.2 13.1 46.0 33.9 35.1
Hymba 16.5 64.0 44.8 36.9 40.0 77.0 53.3 63.1 20.5 59.0 44.2 43.1
   w/ Resona 29.3 69.0 48.4 45.4 51.7 82.0 59.3 68.9 20.9 60.0 45.2 44.0
Table 3: Results on QA benchmarks, where augmentation with Resona improves performance over all evaluation metrics. Hymba indicates Resona is added as a third branch with both the Hymba’s original linear and attention branches, while Hymba(256) indicates where we reduce the window size of Hymba SWA from 1024 to 256, in order to create a fair comparison with our chunk size of 256.
5 Analysis and Disscussion
5.1 Resona  VS RAG-like methods.
A natural comparison to make with Resona is typical RAG-like methods, which append the retrieved passages directly to the input prior to applying the model. To compare with such methods, we design the following methods with Resona: 1) Using the input context directly as the data-store from which a pre-trained retriever can retrieve passages directly that are used as context (Retrieval) and 2) we remove the Knowledge Integration component and instead directly append all (decoded) retrieved passages from the Chunk-and-Search procedure to the initial input (RAG). Table 4 shows these results on a Hymba model, again on the tested QA datasets. The Resona-augmented variant remains significantly more performant than the other alternatives, demonstrating the benefits that acting directly on the representations can have.
Model TriviaQA CoQA NarrativeQA
BLEU ROUGE-L METEOR F1 BLEU ROUGE-L METEOR F1 BLEU ROUGE-L METEOR F1
Hymba 15.3 58.0 41.4 34.3 31.8 62.0 42.1 50.0 10.7 39.0 29.1 28.4
@@ 5 Analysis and Disscussion
   w/Resona 25.6 61.0 43.4 41.5 36.5 69.0 48.8 57.2 13.1 46.0 33.9 35.1
Table 4: Performance comparison of Hymba (256 window size), its Resona-enhanced variants and two different RAG-style input.
5.2 Relationship with Hybrid Methods
Some previous works have suggested methods that enable linear recurrent models to improve their in-context learning abilities. Park et al. (2024) introduce a MambaFormer architecture which interleaves self-attention and Mamba layers, improving the ability to learn in-context on tasks in which pure Mamba models struggle. Similarly, Resona can be interpreted as a form of hybrid mixture of attention and recurrence, similar to Dong et al. (2025), with the difference lying in the frequency of attention and its sparsity within different layers. To better compare these two methods, we provide an ablation (Table 5) where we replace Hymba layers with Resona.
@@ 5.2 Relationship with Hybrid Methods
   w/Resona 31.2 57.0 39.2 44.5
In this setting, 3 consecutive layers are trained in either setup. We observe a meaningful increase in performance on CoQA, indicating that for such types of tasks, the mechanism presented by Resona could be more robust and suitable on a number of real world tasks.
5.3 Ablating on the position of layers.
By its nature, a natural question that emerges relates to the ease and effectiveness of determining the layers at which Resona augmentations need to be made. The same results (Table 5) show that the placement of the 3 Resona layers do not have a significant impact on the performance improvement relative to the baseline, which can be improved upon in nearly all ways in which the layers are selected. This highlights a level of robustness of the framework and method, hinting towards an ability to be used for a variety of additional tasks.
@@ 6 Conclusion
Appendix A Details
A.1 Model Definitions
Let $\{(k_{t},v_{t})\}_{t=1}^{L}$ denote a sequence of key-value pairs with $k_{t},v_{t}\in\mathbb{R}^{d}$, and let $(k^{*},v^{*})$ be the target pair. We analyze three components:
  • Linear Recurrence: The hidden state $S_{t}\in\mathbb{R}^{d\times d}$ in a linear recurrent model evolves as:
    $S_{t}=S_{t-1}+v_{t}k_{t}^{\top},\quad S_{0}=\mathbf{0},$ (1)
    with output $o_{\text{linear}}=S_{t}q=\sum_{i=1}^{t}(k_{i}^{\top}q)v_{i}$.
  • Semantic Search: A sentence transformer $f_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{m}$ retrieves top-$C$ keys $K_{\text{ret}}=\{k_{j}\}_{j=1}^{C}$ via cosine similarity:
    $\cos(f_{\theta}(q),f_{\theta}(k_{j}))=\frac{f_{\theta}(q)^{\top}f_{\theta}(k_{j})}{\|f_{\theta}(q)\|_{2}\|f_{\theta}(k_{j})\|_{2}}.$ (2)
  • Hybrid Model: Combines outputs via:
    $o_{\text{hybrid}}=\lambda o_{\text{linear}}+(1-\lambda)o_{\text{attn}},\quad\lambda\in[0,1],$ (3)
    where $o_{\text{attn}}=\sum_{j=1}^{C}\frac{\exp(q^{\top}k_{j})}{\sum_{m=1}^{C}\exp(q^{\top}k_{m})}v_{j}$.
Appendix B Theoretical Background
B.1 Rank Limitation of Linear Recurrence
Lemma B.1 (Rank Bound).
For $L>d$, the hidden state $S_{L}=\sum_{t=1}^{L}v_{t}k_{t}^{\top}$ satisfies:
$\mathrm{rank}(S_{L})\leq d.$ (4)
Proof.
Each term $v_{t}k_{t}^{\top}$ is rank-1. The sum of $L$ rank-1 matrices in $\mathbb{R}^{d\times d}$ cannot exceed rank $d$. ∎
Corollary B.2 (Interference Error).
The linear model’s retrieval error grows as:
$\mathcal{E}_{\text{linear}}=\left\|o_{\text{linear}}-v^{*}\right\|_{2}\leq\|q\|_{2}\sum_{i\neq*}\|k_{i}\|_{2}\|v_{i}\|_{2}.$ (5)
Proof.
From $o_{\text{linear}}=\sum_{i=1}^{L}(k_{i}^{\top}q)v_{i}$, apply Cauchy-Schwarz: $\left\|\sum_{i\neq*}(k_{i}^{\top}q)v_{i}\right\|_{2}\leq\|q\|_{2}\sum_{i\neq*}\|k_{i}\|_{2}\|v_{i}\|_{2}$. ∎
Appendix C Additional Results
C.1 Detailed Pretraining Results
Param PPL ARC Challenge ARC Easy HellaSwag OpenBookQA PIQA PubMedQA RACE Winogrande AVG
Acc Acc Norm Acc Acc Norm Acc Acc Norm Acc Acc Norm Acc Acc Norm Acc Acc Acc Acc
@@ C.1 Detailed Pretraining Results
DeltaNet (same param) 142M 12.946 0.2159 0.256 0.3001 0.2845 0.2592 0.2655 0.178 0.278 0.5305 0.5125 0.34 0.266 0.4728 0.3199
DeltaNet+Resona 145M 12.541 0.2073 0.2449 0.2963 0.2963 0.2686 0.2757 0.156 0.272 0.5419 0.5283 0.334 0.2517 0.5012 0.3210
Table 6: Detailed harness evaluation results from pre-training.
C.2 Detailed Supervised Fine-Tuning Results
@@
Ver. 1 Ver. 2 Ver. 3 Multi-Value Ver. 1 Ver. 2 Ver. 3 Multi-Value Ver. 1 Ver. 2 Ver. 3 Multi-Value
@@
Mamba 20×$\times\mathsf{lr}$ 5.00 3.65 1.80 1.358 5.00 0.85 0.45 0.528 4.90 0.05 0.00 0.669
Mamba 50×$\times\mathsf{lr}$ 5.00 3.00 1.10 1.267 4.95 0.60 0.20 0.533 4.00 0.05 0.00 0.550
@@
Deltanet 20×$\times\mathsf{lr}$ 5.00 4.10 0.60 0.876 5.00 2.60 0.80 1.453 5.00 0.60 0.10 0.973
Deltanet 50×$\times\mathsf{lr}$ 5.00 4.40 1.45 1.676 5.00 1.25 0.35 1.312 5.00 0.45 0.05 0.885
@@
GLA 20×$\times\mathsf{lr}$ 2.55 3.90 0.10 1.310 0.30 0.50 0.00 0.930 0.00 0.10 0.00 0.578
GLA 50×$\times\mathsf{lr}$ 3.00 3.85 0.15 1.353 0.20 0.60 0.00 0.900 0.00 0.05 0.00 0.657
ARC Challenge ARC Easy HellaSwag OpenBookQA PIQA PubMedQA RACE Winogrande AVG
Acc Acc Norm Acc Acc Norm Acc Acc Norm Acc Acc Norm Acc Acc Norm Acc Acc Acc Acc
@@
Mamba+Resona 0.3823 0.3951 0.6907 0.6904 0.4804 0.6292 0.300 0.406 0.7372 0.7383 0.690 0.3445 0.6338 0.5475
Table 8: Detailed harness evaluation results from supervised fine-tuning of different pre-trained LRMs. After undergoing the same SFT as the backbone models, Resona-enhanced models achieve comparable or superior zero-shot lm-harness evaluation scores to baselines. Combined with Table 7, these results demonstrate that the Resonamodule enhances the backbone’s in-context learning (ICL) capability while maintaining its foundational language modeling performance.
C.3 Training Details for Synthetic Benchmarks

### v3

@@ 3.2 Training and Inference
When training Resona, an important consideration is the auto-regressive nature of the model that therefore requires a causal mask. Specific to our implementation, to maintain this during the parallel nature of training, we ensure that for the query at index $i$, only chunks that solely contain information prior to this position in the input are considered within the Chunk-and-Search process. This ensures that the mask ${\bm{M}}$ allows no information from a given token to affect the representation of those ahead of it in the sequence.
@@ 4.2 Main Results
Model Comp. ICR Noisy ICR Fuzzy ICR SC Mem. Average
Transformer 53.4 96.7 99.7 69.4 98.7 89.4 84.1
Mamba 38.3 76.7 74.9 9.3 33.2 88.5 53.5
+Resona 38.2 ($\downarrow$ 0.1) 99.9 ($\uparrow$ 23.2) 100.0 ($\uparrow$ 25.1) 63.4 ($\uparrow$ 54.1) 42.7 ($\uparrow$ 9.5) 88.8 ($\uparrow$ 0.3) 72.1 ($\uparrow$ 18.6)
Mamba2 43.6 96.4 96.7 21.1 93.3 86.9 73.0
+Resona 46.6 ($\uparrow$ 3.0) 100.0 ($\uparrow$ 3.6) 100.0 ($\uparrow$ 3.3) 62.9 ($\uparrow$ 41.8) 93.6 ($\uparrow$ 0.3) 88.1 ($\uparrow$ 1.2) 81.9 ($\uparrow$ 8.9)
RWKV5 36.8 96.4 96.6 12.1 52.7 55.0 58.3
+Resona 40.4 ($\uparrow$ 3.6) 99.7 ($\uparrow$ 3.3) 99.8 ($\uparrow$ 3.2) 59.7 ($\uparrow$ 47.6) 58.0 ($\uparrow$ 5.3) 70.6 ($\uparrow$ 15.6) 71.5 ($\uparrow$ 13.2)
Hyena 42.2 79.3 77.4 9.96 72.8 88.9 61.7
+Resona 42.6 ($\uparrow$ 0.4) 99.9 ($\uparrow$ 20.6) 99.9 ($\uparrow$ 22.5) 66.2 ($\uparrow$ 56.2) 74.3 ($\uparrow$ 1.5) 89.0 ($\uparrow$ 0.1) 78.7 ($\uparrow$ 17.0)
@@
Figure 5 and Table 1 demonstrate that Resona augmentations improves performance across all baselines, some by wide margins. Baseline models are often able to perfectly solve MQAR for shorter sequence lengths or a smaller number of KV-pairs, but fail catastrophically upon increasing either of these values. Resonameanwhile retains near perfect accuracy even after these values. Similarly, models can struggle at specific tasks within the MAD suite, but Resona achieves a significant improvement in performance. On tasks in which the base LRM models show strong performance, no degradation is observed through the addition of Resona, highlighting its flexibility. We observe that this is consistent across multiple models (Poli et al., 2023; Fu et al., 2023), which are incapable of learning on some of the simpler settings but upon the addition of Resona layers are able of maintaining nearly perfect accuracy for arbitrarily long sequences, showcasing its specific benefits in context-recall intensive settings.
@@
DeltaNet (SP) 142M 12.946
+Resona 145M 12.541
@@
RetNet (SP) 146M 15.431
+Resona 142M 14.742
@@
Mamba (SP) 157M 16.173
+Resona 154M 15.943
@@ 4.2.2 Language Modeling
To assess language modeling, we compare a baseline models with one augmented with Resona layers. Here, we train on the WikiText-103 dataset (Merity et al., 2016). In order to train models augmented with Resona, we also augment the dataset. Specifically, we first consolidate all samples from the same Wikipedia entry into single sample, eliminating excessively short title lines or empty lines. We then use a LLaMA3.1-70B (Grattafiori et al., 2024) model to augment each sample such that we can make use of the Resona retrieval mechanism. We then conduct the Chunk-and-Search process offline to create masks, in order to save computation during training. To account for the additional parameters introduced by Resona, we present results in the baseline settings for two version, one where each LRM layer matches exactly that of the augmented model as well as a version where the hidden size of the layer has been increased to match the parameter count of the augmented counterpart. Results in Table 2 demonstrates that integrating Resona consistently achieves lower perplexity than baseline counterparts and their parameter-aligned variants, highlighting its applicability for language modeling. Notably, modifying Hymba’s sliding window mask to our retrieval-based mask significantly improves performance. Furthermore, results on short-context tasks (Table 8) demonstrate no performance degradation, suggesting the ability to maintain performance on tasks that are not recall intensive.
@@ 4.2.3 Direct Fine-Tuning
Table 3: Results on QA benchmarks, where augmentation with Resona improves performance over all evaluation metrics. Hymba indicates Resona is added as a third branch with both the Hymba’s original linear and attention branches, while Hymba(256) indicates where we reduce the window size of Hymba’s sliding window attention from 1024 to 256, in order to create a fair comparison with our chunk size of 256.
@@
+Resona 25.6 61.0 43.4 41.5 36.5 69.0 48.8 57.2 13.1 46.0 33.9 35.1
Hymba 16.5 64.0 44.8 36.9 40.0 77.0 53.3 63.1 20.5 59.0 44.2 43.1
+Resona 29.3 69.0 48.4 45.4 51.7 82.0 59.3 68.9 20.9 60.0 45.2 44.0
5 Analysis and Discussion
Resona  vs. RAG-like methods.
A natural comparison to make with Resona is typical RAG-like methods, which append the retrieved passages directly to the input prior to applying the model. To compare with such methods, we design the following methods with Resona: 1) Using the input context directly as the data-store from which a pre-trained retriever can retrieve passages directly that are used as context (Retrieval) and 2) we remove the Knowledge Integration component and instead directly append all (decoded) retrieved passages from the Chunk-and-Search procedure to the initial input (RAG).
@@ 5 Analysis and Discussion
+Resona 31.2 57.0 39.2 44.5
Rows 5 and 6 in Table 3 shows these results on a Hymba model, again on the tested QA datasets. The Resona-augmented variant remains significantly more performant than the other alternatives, demonstrating the benefits that acting directly on the representations can have.
Relationship with Hybrid Methods.
Some previous works have suggested methods that enable linear recurrent models to improve their in-context learning abilities. Park et al. (2024) introduce a MambaFormer architecture which interleaves self-attention and Mamba layers, improving the ability to learn in-context on tasks in which pure Mamba models struggle. Similarly, Resona can be interpreted as a form of hybrid mixture of attention and recurrence, similar to Dong et al. (2025), with the difference lying in the frequency of attention and its sparsity within different layers. To better compare these two methods, we provide an ablation (Table 4) where we replace Hymba layers with Resona. In this setting, 3 consecutive layers are trained in either setup. We observe a meaningful increase in performance on CoQA, indicating that for such types of tasks, the mechanism presented by Resona could be more robust and suitable on a number of real world tasks.
Ablating on the position of layers.
Due to its nature, a natural question that emerges relates to the ease and effectiveness of determining the layers at which Resona augmentations need to be made. The same results (Table 4) show that the placement of the 3 Resona layers do not have a significant impact on the performance improvement relative to the baseline, which can be improved upon in nearly all ways in which the layers are selected. This highlights a level of robustness of the framework and method, hinting towards an ability to be used for a variety of additional tasks.
Efficiency Evaluation.
Given the architectural modifications that following from the Resona augmentations, a comparison between the efficiency tradeoffs is also necessary. We provide ablations using Mamba, DeltaNet and GLA, comparing both base models and those augmented with Resona to demonstrate these tradeoffs in Table 5-7. While the Resona augmentations do lead to an increase in each factor, the plug-in remains lightweight and does not add significant overhead in computation, particularly for longer sequences. Furthermore, direct comparison with a Transformer shows this method to be significantly more lightweight while retaining Transformer-like performance on our tasks.
Pre-filling Length Transformer Mamba Mamba + Resona DeltaNet DeltaNet + Resona GLA GLA + Resona
2k 29 45 52 53 64 37 46
4k 34 77 88 59 78 43 64
8k 71 149 170 72 103 62 104
16k 173 294 349 106 181 109 202
32k 503 571 653 208 338 202 386
64k 1665 1118 1285 412 652 407 757
128k 6094 2257 2412 807 1289 806 1518
Table 5: The time (in milliseconds) used to pre-fill a context of a pre-specified length. Numbers are rounded to the nearest millisecond.
Pre-filling Length Transformer Mamba Mamba + Resona DeltaNet DeltaNet + Resona GLA GLA + Resona
2k 2679 2972 3329 2945 3686 2749 3528
4k 2758 3044 3360 3042 3772 2777 3543
8k 2866 3155 3499 3023 3789 2774 3523
16k 3389 3164 3613 3057 3829 2869 3689
32k 5759 3491 3910 3080 4207 2912 4044
64k 11144 4119 4656 3171 4404 3145 4352
128k 24050 4747 5601 3611 5278 3509 5076
Table 6: The time (in milliseconds) taken to generate 128 tokens following a prespecified pre-filling length. Numbers are rounded to the nearest millisecond.
Pre-filling Length Transformer Mamba Mamba + Resona DeltaNet DeltaNet + Resona GLA GLA + Resona
2k 3.1 2.8 4.0 2.9 4.8 2.9 3.6
4k 3.5 2.8 4.1 3.0 4.9 3.0 3.7
8k 4.5 3.1 4.4 3.3 5.2 3.2 3.9
16k 6.4 3.7 5.1 3.7 5.5 3.6 4.4
32k 10.2 4.9 6.5 4.6 6.6 4.3 5.4
64k 17.7 7.2 9.2 6.4 8.4 5.9 7.5
128k 32.9 11.9 14.7 10.2 12.3 9.1 11.7
Table 7: The total memory consumption (in GB) given a pre-specified pre-filling length.
@@ 6 Conclusion
Appendix A Additional Results
A.1 Detailed Pretraining Results
Table 8: Detailed lm-evaluation-harness evaluation results from pre-training.
Param Wiki. ARC-C ARC-E Hella. OBQA PIQA PM RACE Wino. AVG
ppl acc acc_n acc acc_n acc acc_n acc acc_n acc acc_n acc acc acc acc
@@ A.1 Detailed Pretraining Results
+(SP) 142M 12.946 0.2159 0.256 0.3001 0.2845 0.2592 0.2655 0.178 0.278 0.5305 0.5125 0.34 0.266 0.4728 0.3199
+Resona 145M 12.541 0.2073 0.2449 0.2963 0.2963 0.2686 0.2757 0.156 0.272 0.5419 0.5283 0.334 0.2517 0.5012 0.3210
A.2 Detailed Supervised Fine-Tuning Results
@@
V1 V2 V3 MV V1 V2 V3 MV V1 V2 V3 MV
@@
20×$\times\mathsf{lr}$ 5.00 3.65 1.80 1.358 5.00 0.85 0.45 0.528 4.90 0.05 0.00 0.669
50×$\times\mathsf{lr}$ 5.00 3.00 1.10 1.267 4.95 0.60 0.20 0.533 4.00 0.05 0.00 0.550
@@
20×$\times\mathsf{lr}$ 5.00 4.10 0.60 0.876 5.00 2.60 0.80 1.453 5.00 0.60 0.10 0.973
50×$\times\mathsf{lr}$ 5.00 4.40 1.45 1.676 5.00 1.25 0.35 1.312 5.00 0.45 0.05 0.885
@@
20×$\times\mathsf{lr}$ 2.55 3.90 0.10 1.310 0.30 0.50 0.00 0.930 0.00 0.10 0.00 0.578
50×$\times\mathsf{lr}$ 3.00 3.85 0.15 1.353 0.20 0.60 0.00 0.900 0.00 0.05 0.00 0.657
Table 10: Detailed lm-evaluation-harness evaluation results from supervised fine-tuning of different pre-trained LRMs. After undergoing the same SFT as the backbone models, Resona-enhanced models achieve comparable or superior zero-shot lm-harness evaluation scores to baselines. Combined with Table 9, these results demonstrate that the Resonamodule enhances the backbone’s in-context learning (ICL) capability while maintaining its foundational language modeling performance.
ARC-C ARC-E Hella. OBQA PIQA PM RACE Wino. AVG
acc acc_n acc acc_n acc acc_n acc acc_n acc acc_n acc acc acc acc
@@
+Resona 0.3823 0.3951 0.6907 0.6904 0.4804 0.6292 0.300 0.406 0.7372 0.7383 0.690 0.3445 0.6338 0.5475
A.3 Training Details for Synthetic Benchmarks